# All for law and law for all: Adaptive RAG Pipeline for Legal Research

### v1

@@ Abstract
[…]
@@
Figarri Keisha1, Prince Singh1, Pallavi1, Dion Fernandes1, Aravindh Manivannan1, Ilham Wicaksono1, Faisal Ahmad1 1Department of Computer Science, University College London
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated strong generative capabilities but often suffer from hallucinations, which is especially detrimental in high-stakes fields like law, where factual inaccuracies cause significant financial and reputational damage. Retrieval-Augmented Generation (RAG) mitigates this by grounding responses in domain-specific documents, providing concrete sources for the LLM to reference. Some studies on improving information retrieval, such as LegalBenchRAG Pipitone and Alami (2024), providing datasets and benchmarks, which we are leveraging in this study to produce an innovative end-to-end RAG pipeline. Building on the foundations laid by LegalBenchRAG, we attempt to explore and optimise pipelines for a RAG-based legal assistant. The goal was to assess the user’s expertise level using sentiment analysis and tailor a response of the appropriate level, referencing the legal documents provided to the RAG. To achieve this, we concentrated our efforts on three main components of the pipeline: Query Translation, Information Retrieval and Response Generation. The focus of each area/team was to optimise the processes for their respective focus, exploring aspects such as chunking methods, embeddings, query translation, evaluation metrics, and LLMs selection and tuning.
Building on this design, our contributions extend beyond prior work by systematically enhancing each stage of the pipeline and demonstrating how open-source methods can rival proprietary systems. Specifically:
@@ 2 Background
Some research collated state-of-the-art (SoTA) RAG methods with benchmarks such as BERGEN Rau et al. (2024), a Python library for straightforward, reproducible end-to-end RAG experiments. This paper proposed and benchmarked SoTA information retrieval methods such as dense-encoder RetroMAE and SPLADE, a sparse model, both of which we considered for our experimental setup.
Due to resource constraints, we are unable to replicate the experiments using OpenAI’s proprietary embedding model. However, from a practical standpoint, adopting high-performing open-source embedding models is more desirable. Therefore, we focus on evaluating multiple open-source models to determine whether any can match or exceed the retrieval performance of LegalBenchRAG. Additionally, while LegalBenchRAG provides strong retrieval benchmarks, it does not explore end-to-end response generation. This leaves an open question: How well does their retrieval method integrate into a complete RAG pipeline incorporating improved query translation and response generation to produce more accurate responses? To investigate this, we addressed three hypotheses:
@@ 3.1 Dataset
We are using the LegalBenchRAG corpus, which was derived from the LegalBench Guha et al. (2023) dataset, a collaboratively constructed reasoning benchmark consisting of 162 tasks covering six different types of legal reasoning. Legal experts annotated queries by highlighting relevant text in source documents. LegalBench-RAG data has two primary components: the original corpus and the QA pairs. The corpus includes .txt documents from our four legal domains. It excludes documents that were not requested/targeted by at least one query in LegalBenchRAG. The QA pairs are directly linked to the documents within the corpus. Each query is associated with a list of relevant snippets from source documents that directly answer the query. For each snippet, the file path, the exact quote, and the precise character indices within the document are provided, ensuring a clear reference to the source, enabling accurate evaluation of retrieval performance.
@@ 4.1 Query Translation
We observed that many queries in the LegalBench dataset follow a rigid pattern, beginning with “Consider …; …," where the text before the semicolon specifies a document reference. To address this, we built a Simple Extractor (SE) that splits queries at the semicolon, removes stopwords (e.g., "Non-Disclosure Agreement"), and embeds the document reference with a sentence transformer (all-MiniLM-L6-v2) to find the best-matching file via cosine similarity. Matches are scored as 1 for correct, -1 for incorrect, and 0 if the similarity falls below a threshold. Because CUAD file names often diverge from the textual content of queries, we increased the threshold to 0.55 for CUAD to reduce mismatches, while lower thresholds (0.3–0.38) were sufficient for the other datasets. Table 2 reports SE’s performance on the original dataset queries with their respective thresholds.
@@
Although SE performs well when queries strictly follow the "Consider …; …" pattern, many queries deviate from this format in real usage. To test robustness, we rephrased all queries in the datasets using a few-shot prompt, ensuring that the document reference (e.g., "In the Non-Disclosure Agreement between CopAcc and ToP Mentors …") was preserved while rewriting the question into a more conversational style (e.g., "Is it clearly stated that the Receiving Party has no rights …?"). We employed Flan-T5 Large to generate these more natural queries, approximating real-world scenarios.
@@
  1. 1.
    Expertise Classification: We employed the Dale-Chall readability formula to estimate the linguistic complexity of each query for all 4 datasets. It was chosen for its proven ability to assess comprehensibility in formal domains like legal text. Unlike metrics based only on sentence length or syllables, it uses a curated list of familiar words, making it better suited for structured, jargon-heavy text, an observation supported by Han et al. (2024) for legal texts and Zheng and Yu (2017) for technical health documents. To distinguish between domain experts and laypersons, we used a threshold: scores below 8.0 were labelled non-expert, and 8.0 or above as expert, thus guiding the response generation module to produce answers that are either technically detailed or simplified for broader accessibility.
  2. 2.
    Vague vs Verbose Classification: To adapt the retrieval strategy based on query specificity, we introduced a classification mechanism that labels queries as vague or verbose. This distinction enables the system to dynamically adjust the number of chunks retrieved during grounding: vague queries, which are general or under-specified, benefit from broader retrieval, while verbose queries, often multi-part or over-specified, require more selective, targeted context windows. Following HyPA-RAG Kalra et al. (2025), we constructed a synthetic dataset by rephrasing LegalBench queries for all 4 datasets with Meta-LLaMA-3 to create diverse vague and verbose variants. This dataset was used to train DistilBERT for binary classification. Once classified, the vague or verbose label was passed to the retrieval module, guiding chunking behaviour and context scaling for downstream LLM processing.
@@ 4.2 Information Retrieval
  1. 1.
    Embedding Model: We compared three open-source embedding models: SBERT (all-mpnet-base-v2), Distilled SBERT (all-MiniLM-L6-v2), and GTE-Large (thenlper/gte-large) against LegalBenchRAG’s text-embedding-3-large. These models span a range of scales: mpnet-base (110M) offers high accuracy, MiniLM (22M) is lightweight and efficient, and GTE-Large is optimised for retrieval and reranking in RAG pipelines. This selection allowed us to evaluate how model size and design influence retrieval performance.
  2. 2.
    Pipeline: The sampled corpus was chunked using Naïve and RCTS methods, embedded with three models, and stored as JSON vectors. Queries from the benchmarks were also embedded, and similarity search (cosine and BM25) retrieved the top 50 chunks, referred to as unranked for comparison purposes with reranked chunks. However, they are ordered by similarity scores. A reranker then reordered these based on query similarity to produce reranked chunks. Precision and recall were calculated by comparing unranked and reranked results against ground truth spans across chunking–embedding–similarity combinations and k-values (1–50). Additionally, we tested RetroMAE, a SoTA embedding model from the BERGEN paper Rau et al. (2024), against the best-performing configuration to evaluate its effectiveness in the RAG pipeline.
  3. 3.
    Evaluation Approach: Following LegalBenchRAG, we used Precision@K and Recall@K based on span overlap. Retrieved chunks were compared to benchmark QA ground truth by file and span alignment. Precision was computed as overlap length over chunk length, and recall as overlap over ground truth length. With multiple chunks and answers per query, scores were averaged across 194 QA pairs per domain and overall, over varying K-values (see Figure 2). After identifying the best model, we extended the evaluation to include a SoTA method and increased K to 300 for direct comparison with LegalBenchRAG. This extended evaluation was run only on the top-performing model due to resource constraints.
    We explored using text-based evaluation in addition to span-based evaluation as a way to measure sentiment similarity between retrieved chunks and ground truth. Ultimately, we decided to proceed using primarily span-based, and used the text-based evaluation as a sanity check by paying attention to the behaviour of precision and recall across K-values, such as in Figure 6 in Appendix B.
@@ 4.3 Response Generation and Evaluation (RGE)
  1. 1.
    Prompt Designs: We experimented with several prompts: i) baseline is the RAG prompt from LangChain, ii) zero-shot Chain of Thought (CoT) prompt to assess potential reasoning improvements, and finally iii) a custom-crafted prompt with explicit instructions to enhance accuracy, legal grounding, and relevance.
  2. 2.
    Evaluation Metrics: RAGAS (RAG Assessment) is used, especially Answer Relevancy and Faithfulness Exploding Gradients (2025). The original RAGAS paper proposed Context Relevance, but it was later deemed unhelpful by the authors Es et al. (2023). Instead, we focus on two RAGAS reference-free metrics: Answer Relevancy, which compares LLM-generated questions (from the model’s response) to the original question using similarity scores, and Faithfulness, which checks if the retrieved context supports LLM-generated claims from the response. In addition, we consider using two reference-based metrics: BERTScore-F1 (with LegalBERT) to measure semantic similarity between generated answers and contexts, and ROUGE-Recall to assess completeness through n-gram overlap, to support faithfulness assessment.
  3. 3.
    Models: GPT-4o-mini, and LLAMA-3-8B.
  4. 4.
    Response generation framework: Before fitting to the final pipeline and using query translation outputs, we want to find the optimal combination of prompt, model, and k-value by generating responses for all possible combinations using the chunks retrieved by the best model. K-values were varied (1, 3, 5, 10) to compare its effect on generation performance.
@@ 6.1 Recall@K and Precision@K
We observed a few similar findings with LegalBenchRAG. Firstly, RCTS performs better than Naive chunking, although not consistently for all model combinations. Some experiments, such as Naive+SBERT+Cosine and Naive+GTE+Cosine, were in the top 5 performers, at times performing well for individual domains. However, considering overall performance, the top 2 model combinations used RCTS chunking. Thus, overall we can conclude that RCTS outperformed Naive chunking. Secondly, unranked results perform better than reranked ones for cosine similarity and conversely for BM25 similarity, proving our Hypothesis 3 inconclusive and dependent on model choice. LegalBenchRAG also observed that unranked performs better with cosine similarity. Figure 2 shows these comparisons. Figure 2 also shows that cosine similarity performs better than BM25 similarity search, answering Hypothesis 2. The performances may vary for each domain, which is presented in Figure 7 in Appendix C.
@@ 6.2 Response Generation and Evaluation
In considering the optimal K-value, we note that the Faithfulness stagnates after K=5 and BERT-F1 stagnates with marginal fluctuations for all K-values. Based on Figure 2, we can see that having too high a K-value will result in a loss of precision but a gain on recall for information retrieval, thus requiring a fine balance in choosing the optimal K-value to trade-off precision and recall. Thus, we can consider a K around 5, which has good Faithfulness, BERT-F1, precision, and decent recall.
For the second part of the evaluation, which incorporates the Readability and Complexity classifier (R&C), we applied K==5 for non-expert queries and increased K to 10 for expert queries, to add more details for complex questions. Figure 5 shows that including R&C does not change the performance significantly.
@@
7 Conclusions
LLMs often suffer from hallucinations, which is a critical issue in the legal domain where information is crucial. This study addresses this key issue by successfully optimising an end-to-end RAG pipeline for legal documents, advancing beyond the LegalBenchRAG by integrating query translation, retrieval and response generation. Using open-source models (SBERT, GTE-large) enabled cost-effective and more accessible pipelines without the reliance on commercial APIs. The pipeline incorporates a novel strategy including query translation and complexity-driven chunking and generation. Complexity classification (i.e. expert vs. non-expert and vague vs. verbose) has helped to reduce hallucinations and tailor retrieval. The simple extractor used on structured queries helped to significantly improve the recall and precision for retrieval. For the retrieval stage, RCTS and cosine similarity outperformed Naïve chunking and BM25. Query translation significantly increased recall@K and Precision@K for information retrieval, while adaptive top-k retrieval based on complexity outperforms fixed-k in response generation. Furthermore, faithfulness and BERT-F1 evaluated legal response quality better than ROUGE. This highlights the value of adapting retrieval and generation according to query complexity, providing a strong foundation for lightweight legal tools without commercial APIs.

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) has transformed how we approach text generation tasks by grounding Large Language Model (LLM) outputs in retrieved knowledge. This capability is especially critical in the legal domain. In this work, we introduce a novel end-to-end RAG pipeline that improves upon previous baselines using three targeted enhancements: (i) a context-aware query translator that disentangles document references from natural-language questions and adapts retrieval depth and response style based on expertise and specificity, (ii) open-source retrieval strategies using SBERT and GTE embeddings that achieve substantial performance gains while remaining cost-efficient, and (iii) a comprehensive evaluation and generation framework that combines RAGAS, BERTScore-F1, and ROUGE-Recall to assess semantic alignment and faithfulness across models and prompt designs. Our results show that carefully designed open-source pipelines can rival proprietary approaches in retrieval quality, while a custom legal-grounded prompt consistently produces more faithful and contextually relevant answers than baseline prompting. Taken together, these contributions demonstrate the potential of task-aware, component-level tuning to deliver legally grounded, reproducible, and cost-effective RAG systems for legal research assistance.
@@
Figarri Keisha1, Prince Singh1, Pallavi1, Dion Fernandes1, Aravindh Manivannan1, Ilham Wicaksono1, Faisal Ahmad1, Wiem Ben Rim1 1 University College London
@@ 1 Introduction
Large Language Models demonstrate strong generative capabilities that make them incredibly versatile and useful in various domains and applications. However, they remain vulnerable to hallucinations, which proves to be especially detrimental in high-stakes fields (e.g law), where factual inaccuracies cause significant financial and reputational damage. Retrieval-Augmented Generation (RAG) mitigates this by grounding responses in domain-specific documents, providing concrete sources for the LLM to reference. Previous work investigates the improvement of RAG models for the legal domain, for instance LegalBenchRAG Pipitone and Alami (2024) introduces new datasets and benchmarks to better evaluate models which we leverage for our pipeline.
In this work, we build and optimize a novel end-to-end RAG pipeline that can adapt to the user’s needs. For this purpose, we introduce three main components of the pipeline: Query Translation, Information Retrieval and Response Generation. First, we assess the user’s expertise level. A layperson asking a question about a social networking collecting their data would be different from a legal professional asking the same question. Once the user expertise is estimated, we tailor our responses while referencing related legal documents in our database. For this purpose, our contributions extend beyond prior work by systematically enhancing each stage of the pipeline and demonstrating how open-source methods can rival proprietary systems. Specifically:
@@ 2 Background
From a practical standpoint, adopting high-performing open-source embedding models is more desirable. Therefore, we focus on evaluating multiple open-source models to determine whether any can match or exceed the retrieval performance of LegalBenchRAG. Additionally, while LegalBenchRAG provides strong retrieval benchmarks, it does not explore end-to-end response generation. This leaves an open question: How well does their retrieval method integrate into a complete RAG pipeline incorporating improved query translation and response generation to produce more accurate responses? To investigate this, we addressed three hypotheses:
@@ 3.1 Dataset
We use the LegalBenchRAG corpus, which was derived from the LegalBench Guha et al. (2023) dataset, a collaboratively constructed reasoning benchmark consisting of 162 tasks covering six different types of legal reasoning. Legal experts annotated queries by highlighting relevant text in source documents. LegalBench-RAG data has two primary components: the original corpus and the QA pairs. The corpus includes .txt documents from four legal domains: ContractNLI (contract understanding and natural language inference), CUAD (Contract Understanding Atticus Dataset for reviewing legal contracts), MAUD (merger and acquisition understanding dataset), and PrivacyQA (privacy policy question answering). It excludes documents that were not requested/targeted by at least one query in LegalBenchRAG. The QA pairs are directly linked to the documents within the corpus. Each query is associated with a list of relevant snippets from source documents that directly answer the query. For each snippet, the file path, the exact quote, and the precise character indices within the document are provided, ensuring a clear reference to the source, enabling accurate evaluation of retrieval performance.
@@ 4.1 Query Translation
We observe that many entries in the LegalBench dataset follow a rigid pattern specifically mentioning a reference document along with a query in the format “Consider document A; query B“. To address this, we build a Simple Extractor (SE) that processes the queries by splitting them, removing stopwords (e.g., "Non-Disclosure Agreement"), and embeding the document reference with a sentence transformer (all-MiniLM-L6-v2) to find the best-matching file via cosine similarity. Matches are scored as 1 for correct, -1 for incorrect, and 0 if the similarity falls below a threshold. With the exception of CUAD (Contract Understanding Atticus Dataset) where file names often diverge from the textual content of queries due to their standardized naming convention, we increase the threshold to 0.55 for CUAD to reduce mismatches, while lower thresholds (0.3–0.38) are sufficient for the other datasets. Table 2 reports SE’s performance on the original dataset queries with their respective thresholds.
@@
Although SE performs well when queries strictly follow aforementioned rigid pattern, many queries deviate from this format in real usage. To test robustness, we rephrase all queries in the datasets using a few-shot prompt, ensuring that the document reference (e.g., "In the Non-Disclosure Agreement between CopAcc and ToP Mentors …") is preserved while rewriting the question into a more conversational style (e.g., "Is it clearly stated that the Receiving Party has no rights …?"). We employ Flan-T5 Large to generate these more natural queries, approximating real-world scenarios.
@@
Expertise Classification:
We employ the Dale-Chall readability formula to estimate the linguistic complexity of each query for all 4 datasets. It is chosen for its proven ability to assess comprehensibility in formal domains like legal text. Unlike metrics based only on sentence length or syllables, it uses a curated list of familiar words, making it better suited for structured, jargon-heavy text, an observation supported by Han et al. (2024) for legal texts and Zheng and Yu (2017) for technical health documents. To distinguish between domain experts and laypersons, we use a threshold: scores below 8.0 are labelled non-expert, and 8.0 or above as expert, thus guiding the response generation module to produce answers that are either technically detailed or simplified for broader accessibility.
Vague vs Verbose Classification:
To adapt the retrieval strategy based on query specificity, we introduced a classification mechanism that labels queries as vague or verbose. This distinction enables the system to dynamically adjust the number of chunks retrieved during grounding: vague queries, which are general or under-specified, benefit from broader retrieval, while verbose queries, often multi-part or over-specified, require more selective, targeted context windows. Following HyPA-RAG Kalra et al. (2025), we constructed a synthetic dataset by rephrasing LegalBench queries for all 4 datasets with Meta-LLaMA-3 to create diverse vague and verbose variants. This dataset was used to train DistilBERT for binary classification. Once classified, the vague or verbose label was passed to the retrieval module, guiding chunking behaviour and context scaling for downstream LLM processing.
@@ 4.2 Information Retrieval
Embedding Model:
We compare three open-source embedding models: SBERT (all-mpnet-base-v2), Distilled SBERT (all-MiniLM-L6-v2), and GTE-Large (thenlper/gte-large) against LegalBenchRAG’s text-embedding-3-large. These models span a range of scales: mpnet-base (110M) offers high accuracy, MiniLM (22M) is lightweight and efficient, and GTE-Large is optimised for retrieval and reranking in RAG pipelines. This selection allows us to evaluate how model size and design influence retrieval performance.
Pipeline:
The sampled corpus was chunked using Naïve and RCTS methods, embedded with three models, and stored as JSON vectors. Queries from the benchmarks were also embedded, and similarity search (cosine and BM25) retrieved the top 50 chunks, referred to as unranked for comparison purposes with reranked chunks. However, they are ordered by similarity scores. A reranker then reordered these based on query similarity to produce reranked chunks. Precision and recall were calculated by comparing unranked and reranked results against ground truth spans across chunking–embedding–similarity combinations and k-values (1–50). Additionally, we tested RetroMAE, a SoTA embedding model from the BERGEN paper Rau et al. (2024), against the best-performing configuration to evaluate its effectiveness in the RAG pipeline.
Evaluation Approach:
Following LegalBenchRAG, we use Precision@K and Recall@K based on span overlap. Retrieved chunks are compared to benchmark QA ground truth by file and span alignment. Precision is computed as overlap length over chunk length, and recall as overlap over ground truth length. With multiple chunks and answers per query, scores are averaged across 194 QA pairs per domain and overall, over varying K-values (see Figure 2). After identifying the best model, we extend the evaluation to include a SoTA method and increase K to 300 for direct comparison with LegalBenchRAG. This extended evaluation is run only on the top-performing model due to resource constraints.
We explore using text-based evaluation in addition to span-based evaluation as a way to measure sentiment similarity between retrieved chunks and ground truth. Ultimately, we decide to proceed using primarily span-based, and use the text-based evaluation as a sanity check by paying attention to the behaviour of precision and recall across K-values, such as in Figure 6 in Appendix B.
@@ 4.3 Response Generation and Evaluation (RGE)
Prompt Designs:
We experiment with several prompts: i) baseline is a standard RAG prompt, ii) zero-shot Chain of Thought (CoT) prompt to assess potential reasoning improvements, and finally iii) a custom-crafted prompt with explicit instructions to enhance accuracy, legal grounding, and relevance.
Evaluation Metrics:
RAGAS (RAG Assessment) is used, especially Answer Relevancy and Faithfulness Exploding Gradients (2025). The original RAGAS paper proposed Context Relevance, but it was later deemed unhelpful by the authors Es et al. (2023). Instead, we focus on two RAGAS reference-free metrics: Answer Relevancy, which compares LLM-generated questions (from the model’s response) to the original question using similarity scores, and Faithfulness, which checks if the retrieved context supports LLM-generated claims from the response. In addition, we consider using two reference-based metrics: BERTScore-F1 (with LegalBERT) to measure semantic similarity between generated answers and contexts, and ROUGE-Recall to assess completeness through n-gram overlap, to support faithfulness assessment.
Models:
GPT-4o-mini, and LLAMA-3-8B.
Response generation framework:
Before fitting to the final pipeline and using query translation outputs, we want to find the optimal combination of prompt, model, and k-value by generating responses for all possible combinations using the chunks retrieved by the best model. K-values were varied (1, 3, 5, 10) to compare its effect on generation performance.
@@ 6.1 Retrieval
We observe a few similar findings with LegalBenchRAG. Firstly, RCTS performs better than Naive chunking, although not consistently for all model combinations. Some experiments, such as Naive+SBERT+Cosine and Naive+GTE+Cosine, are in the top 5 performers, at times performing well for individual domains. However, considering overall performance, the top 2 model combinations use RCTS chunking. Thus, overall we can conclude that RCTS outperforms Naive chunking. Secondly, unranked results perform better than reranked ones for cosine similarity and conversely for BM25 similarity, proving our Hypothesis 3 inconclusive and dependent on model choice. LegalBenchRAG also observes that unranked performs better with cosine similarity. Figure 2 shows these comparisons. Figure 2 also shows that cosine similarity performs better than BM25 similarity search, answering Hypothesis 2. The performances may vary for each domain, which is presented in Figure 7 in Appendix C.
@@ 6.2 Response Generation and Evaluation
In considering the optimal K-value, we note that the faithfulness stagnates after K=5 and BERT-F1 stagnates with marginal fluctuations for all K-values. Based on Figure 2, we can see that having too high a K-value will result in a loss of precision but a gain on recall for information retrieval, thus requiring a fine balance in choosing the optimal K-value to trade-off precision and recall. For these reasons, we choose K=5.
For the second part of the evaluation, which incorporates the Readability and Complexity classifier (R&C), we apply K==5 for non-expert queries and increase K to 10 for expert queries, to add more details for complex questions. Figure 5 shows that including R&C does not change the performance significantly.
@@
7 Conclusion
LLMs often suffer from hallucinations, which is a critical issue in the legal domain where obtaining the correct information is crucial. This study addresses this key issue by successfully optimising an end-to-end RAG pipeline for legal documents, advancing beyond the LegalBenchRAG by integrating query translation, retrieval and response generation.
Our findings highlight the value of adapting retrieval and generation according to query complexity, providing a strong foundation for lightweight legal tools without commercial APIs.